# CLCE: An Approach to Refining Cross-Entropy and Contrastive Learning for Optimized Learning Fusion

### v1

@@ 1 Introduction
Approaches for achieving state-of-the-art performance in image classification tasks often employ models initially pre-trained on auxiliary tasks and then fine-tuned on a task-specific labeled dataset with a Cross-Entropy loss (CE) (Dosovitskiy et al., 2021; Wang et al., 2022; He et al., 2022; Long and McCreadie, a, b; Long et al., d; Yi et al., 2023). However, CE’s inherent limitations can impact model performance. Specifically, the measure of KL-divergence between one-hot label vectors and model outputs can cause narrow decision margins in the feature space. This hinders generalization (Liu et al., 2016; Cao et al., 2019) and has been shown to be sensitive to noisy labels (Nar et al., 2019; Liu et al., 2016) or adversarial samples (Elsayed et al., 2018; Nar et al., 2019). Various techniques have emerged to address these problems, such as knowledge distillation (Hinton et al., 2015), self-training (Yalniz et al., 2019), Mixup (Zhang et al., 2017), CutMix (Yun et al., 2019), and label smoothing (Szegedy et al., 2015). However, in scenarios such as few-shot learning, these issues with CE have not been fully mitigated. Indeed, while techniques such as extended fine-tuning epochs and specialized optimizers (Zhang et al., 2020; Mosbach et al., 2020) can reduce the impact of CE to some extent, they introduce new challenges, such as extended training time and increased model complexity (Zhang et al., 2020; Mosbach et al., 2020; Long et al., a).
@@
Amidst these challenges in context of image classification, contrastive learning has emerged as a promising solution (Long et al., b), particularly in few-shot learning scenarios such as CIFAR-FS (Bertinetto et al., 2019) and CUB-200-2011 datasets (Wah et al., 2011). The effectiveness of contrastive learning lies in its ability to amplify similarities among positive pairs (intra-class data points) and distinguish negative pairs (inter-class data points). SimCLR (Chen et al., 2020), for instance, has utilized instance-level comparisons unsupervised. However, this unsupervised approach raises concerns regarding its effectiveness , primarily because it limits the positive pairs to be transformed views of an image and treats all other samples in a mini-batch as negatives, potentially overlooking actual positive pairs. We hypothesis incorporating task-specific label information is thus crucial for accurately identifying all positive pairs, especially given the presence of labels in many downstream datasets.
@@
There is a growing trend of using task labels with contrastive learning to replace the standard use of CE (Khosla et al., 2020). A critical observation here is that many state-of-the-art methods, both in supervised (Khosla et al., 2020; Gunel et al., 2020) and unsupervised (Chen et al., 2020; Sohn, 2016; Ge et al., 2018) contrastive learning, overlook the strategic selection of negative samples. They fail to differentiate or prioritize these samples during selection or processing, thereby missing the benefits of leveraging “hard” negative samples, as highlighted in numerous studies (Song et al., 2015; Kumar et al., 2017; Suh et al., 2019; Chu et al., 2021; Zhao et al., 2021; Li et al., 2022). While contrastive learning mitigates the limitations of CE, it simultaneously introduces a challenge: a reliance on large batch sizes—such as 2048 or 4096 samples per batch—for superior performance compared to CE. This requirement is often impractical in budget hardware environments, particularly when using GPUs with less than 24 GB of memory. As a consequence, state-of-the-art methods such as SupCon (Khosla et al., 2020) underperform compared to CE when using more commonly employed batch sizes, such as 64 or 128 samples per batch, which limits their application. Motivated by these successes and gaps in research, we pose the question: How can the performance of contrastive learning be improved to address the shortcomings of cross-entropy loss, while also mitigating the reliance on large batch sizes?
@@
  • State-of-the-Art Performance in Few-Shot Learning and Transfer Learning settings: CLCE significantly surpasses CE by an average of 2.74% in Top-1 accuracy across four few-shot learning datasets when using the BEiT-3 base model (Wang et al., 2022), with large gains observed in 1-shot learning scenarios. Additionally, in transfer learning settings, CLCE consistently outperforms other state-of-the-art methods across eight image datasets, setting a new state-of-the-art result for base models (88 million parameters) on ImageNet-1k (Deng et al., 2009).
  • Reduced Contrastive Learning’s Dependency on Large Batch Sizes: Empirical evidence shows that CLCE significantly outperforms both CE and previous state-of-the-art contrastive learning methods like SupCon (Khosla et al., 2020) in commonly used batch sizes, such as 64. This is a size at which earlier state-of-the-art contrastive learning methods underperform. This advancement tackles a crucial bottleneck in contrastive learning, particularly in settings with limited resources. It positions CLCE as a viable, efficient alternative to conventional CE.
@@ 2.1 Limitations of Cross-Entropy loss
The cross-entropy loss (CE) has long been the default setting for many deep neural models due to its ability to optimize classification tasks effectively. However, recent research has revealed several inherent drawbacks (Liu et al., 2016; Cao et al., 2019; Long et al., c). Specifically, models trained with the CE tend to exhibit poor generalization capabilities. This vulnerability stems from the model having narrow decision margins in the feature space, making it more susceptible to errors introduced by noisy labels (Nar et al., 2019; Liu et al., 2016) and adversarial examples (Elsayed et al., 2018; Nar et al., 2019). These deficiencies underscore the need for alternatives that offer better robustness and discrimination capabilities.
@@ 2.2 Contrastive Learning and Negative Mining
The exploration of negative samples, particularly hard negatives, in contrastive learning has emerged as a critical yet relatively underexplored area. While the significance of positive sample identification is well-established, recent studies have begun to unravel the intricate role of hard negatives. The potential of hard negative mining in latent spaces has been validated in numerous studies (Suh et al., 2019; Chu et al., 2021; Zhao et al., 2021; Li et al., 2022; Wu et al., 2022; Yang et al., 2022; Long et al., 2023a, b). These studies highlight the pivotal role of hard negatives in enhancing the discriminative capability of embeddings. In the contrastive learning domain, Chuang et al. (2020) tackled the challenge of discerning true negatives from a vast pool of candidates by approximating the true negative distribution. Later, Robinson et al. (2021) applied hard negative mining to unsupervised contrastive learning, resulting in a framework where only a single positive pair is utilized in each iteration of the loss calculation. However, these approaches still presents limitations, such as inaccurately identifying positive and negative samples and only using one positive pair, which harms the performance of contrastive learning. Jiang et al. (2022) expand upon the concept of hard negative mining within a supervised framework. Their approach employs a consistent threshold-based dot product for identifying “hard” samples. However, determining an appropriate threshold remains challenging, as it varies significantly across different datasets and even within individual mini-batches. Moreover, their methodology does not tackle the dependency on large batch sizes, which is a critical limitation on performance and applicability.
@@ 3.1 CLCE
$\mathcal{L}_{\text{LACLN}}=\displaystyle\sum\limits_{x_{i}\in\mathcal{D}^{*}}\log\frac{-1}{\lvert\mathcal{D}^{*\textbf{+}}_{-x_{i}}\lvert}\frac{\displaystyle\sum\limits_{x_{p}\in\mathcal{D}^{*\textbf{+}}_{-x_{i}}}\exp(x_{i}\cdot x_{p}/\tau)}{\displaystyle\sum\limits_{x_{p}\in\mathcal{D}^{*\textbf{+}}_{-x_{i}}}\exp(x_{i}\cdot x_{p}/\tau)+\displaystyle\sum\limits_{x_{k}\in\mathcal{D}^{*\textbf{-}}_{-x_{i}}}\displaystyle\frac{\lvert\mathcal{D}^{*\textbf{-}}_{-x_{i}}\lvert}{\displaystyle\sum\limits_{x_{k}\in\mathcal{D}^{*\textbf{-}}_{-x_{i}}}{\exp(x_{i}\cdot x_{k}/\tau)}}\exp(x_{i}\cdot x_{k}/\tau)^{2}}$
@@ 3.3 Representation Learning Framework
  • Data Augmentation module, ${Aug(\cdot)}$: This component creates two different views of each sample $x$, denoted $\tilde{x}={Aug(x)}$. This means that every sample will have at least one similar sample (positive pair) in a batch during training.
  • Encoder Network, ${Enc(\cdot)}$: This network encodes the input data, $x$, into a representation vector, $r=Enc(x)$. Each of the two different views of the data is fed into the encoder separately.
  • Classification head, $Head(\cdot)$: This maps the representation vector, $r$, to probabilities of classes in the target task. The mapping primarily consists of a linear layer, and we utilize its output to calculate the cross-entropy loss.
Our CLCE approach (Eq. 3) can be applied using a wide range of encoders, such as BEiT-3 (Wang et al., 2022) or the ResNets (He et al., 2016) for image classification. Following the method in Chen et al. (2020), every image in a batch is altered to produce two separate views (anchors). Views with the same label as the anchor are considered positive, while the rest are viewed as negative. The encoder output, represented by $y_{i}={Enc(z_{i})}$, is used to calculate the contrastive loss. In contrast, the output from the classification head, denoted as $p_{i}={Head(Enc(z_{i}))}$, is used for the CE. We have incorporated L2 normalization on encoder outputs, a strategy demonstrated to enhance performance significantly (Tian et al., 2020).
@@ 4 Evaluation
We employ representative models from two categories of architectures – BEiT-3/MAE/ViT base (Wang et al., 2022; He et al., 2022; Dosovitskiy et al., 2021) (transformers based models), and ResNet-101 (He et al., 2016) (convolutional neural network). While new state-of-the-art models are continuously emerging (e.g. DINOv2 (Oquab et al., 2023)), our focus is not on the specific choice of architecture. Instead, we aim to show that CLCE is model-agnostic by demonstrating performance gains with two very different and widely used architectures, as well as show it can be trained and deployed in hardware-constrained settings. Further implementation details and the complete code for all experiments can be found in the supplementary material, which will be made publicly available upon acceptance.
@@
(Dhillon et al., 2020) Transductive 76.58$\pm$0.68 85.79$\pm$0.50 43.16$\pm$0.59 57.57$\pm$0.55 65.73$\pm$0.68 78.40$\pm$0.52 73.34$\pm$0.71 85.50$\pm$0.50
(Zhang et al., 2021) Meta-QDA 75.83$\pm$0.88 88.79$\pm$0.75 - - 67.83$\pm$0.64 84.28$\pm$0.69 74.33$\pm$0.65 89.56$\pm$0.79
(Hiller et al., 2022) FewTRUE-ViT 76.10$\pm$0.88 86.14$\pm$0.64 46.20$\pm$0.79 63.14$\pm$0.73 68.02$\pm$0.88 84.51$\pm$0.53 72.96$\pm$0.92 87.79$\pm$0.67
(Hiller et al., 2022) FewTRUE-Swin 77.76$\pm$0.81 88.90$\pm$0.59 47.68$\pm$0.78 63.81$\pm$0.75 72.40$\pm$0.78 86.38$\pm$0.49 76.32$\pm$0.87 89.96$\pm$0.55
(Hu et al., 2023) BAVARDAGE 82.68$\pm$0.25 89.97$\pm$0.18 52.60$\pm$0.32 65.35$\pm$0.25 77.85$\pm$0.28 88.02$\pm$0.14 79.38$\pm$0.29 88.04$\pm$0.18
@@ 4.1 Few-shot Learning
We evaluate our proposed CLCE in the few-shot learning setting, i.e. each test run comprises 3,000 randomly sampled tasks, and we report median Top-1 accuracy with a 95% confidence interval across three runs, maintaining a consistent query shot count of 15. Four prominent benchmarks are used for evaluation: CIFAR-FS (Bertinetto et al., 2019), FC100 (Oreshkin et al., 2018), miniImageNet (Vinyals et al., 2016), and tieredImageNet (Ren et al., 2018). We follow established splitting protocols for a fair comparison (Bertinetto et al., 2019; Oreshkin et al., 2018; Ravi and Larochelle, 2017).
Tab. 1 shows the performance of BEiT-3 and ResNet-101 models under various methods, including CE, H-SCL (Jiang et al., 2022), and the same weighted combination of CE and state-of-the-art supervised contrastive learning loss (SupCon) (Khosla et al., 2020) as CLCE. The results reveal that our CLCE approach consistently improves classification accuracy over other methods, demonstrating superior generalization with limited training data for each class. Our CLCE enhances models’ performance on few-shot datasets, significantly outperforming both CE and CE+SuperCon (paired t-test, p $<$ 0.01). In the 1-shot learning context when compared to BEiT-3 trained with CE (BEiT-3-CE), the most remarkable improvement is seen on the FC100 dataset, with accuracy rising by $3.52\%$ through the use of CLCE (BEiT-3-CLCE). Indeed, across all datasets, BEiT-3-CLCE shows an average accuracy improvement of $2.7\%$. For 5-shot learning, the average improvements across the datasets are $1.4\%$ in accuracy for BEiT-3-CLCE, demonstrating CLCE’s effectiveness in scenarios with fewer positive samples per class and its ability to yield consistent and reliable results, evident in the tighter confidence intervals for Top-1 accuracy. As for ResNet-101, CLCE (ResNet-101-CLCE) demonstrates even more significant improvements over both CE and CE+SupCon. The enhancement is especially remarkable in the case of tieredImagenet, where ResNet-101-CLCE achieves increases of $16.68\%$ over ResNet-101-CE and $14.17\%$ over ResNet-101-CE+SupCon in 1-shot learning. For 5-shot learning, the improvements are $16.9\%$ and $13.36\%$, respectively. On average, ResNet-101-CLCE achieves a $9.82\%$ improvement in 1-shot and an $8.71\%$ improvement in 5-shot settings over the ResNet-101-CE. Lastly, H-SCL (Jiang et al., 2022) underperforms compared to CE at a batch size of 128. This highlights contrastive learning’s limitation of needing very large batch sizes for better performance than CE, evident in ResNet-101 and BEiT-3 models. Overall, these outcomes underline the efficacy of our proposed CLCE approach and CLCE’s broad applicability across different model architectures for few-shot learning tasks.

### v2

@@ 1 Introduction
Approaches for achieving state-of-the-art performance in image classification tasks often employ models initially pre-trained on auxiliary tasks and then fine-tuned on a task-specific labeled dataset with a Cross-Entropy loss (CE) [9, 56, 17, 28, 29, 31, 63]. However, CE’s inherent limitations can impact model performance. Specifically, the measure of KL-divergence between one-hot label vectors and model outputs can cause narrow decision margins in the feature space. This hinders generalization [27, 3] and has been shown to be sensitive to noisy labels [39, 27] or adversarial samples [10, 39]. Various techniques have emerged to address these problems, such as knowledge distillation [19], self-training [61], Mixup [65], CutMix [64], and label smoothing [51]. However, in scenarios such as few-shot learning, these issues with CE have not been fully mitigated. Indeed, while techniques such as extended fine-tuning epochs and specialized optimizers [66, 37] can reduce the impact of CE to some extent, they introduce new challenges, such as extended training time and increased model complexity [66, 37, 36].
@@
Amidst these challenges in context of image classification, contrastive learning has emerged as a promising solution [33], particularly in few-shot learning scenarios such as CIFAR-FS [2] and CUB-200-2011 datasets [55]. The effectiveness of contrastive learning lies in its ability to amplify similarities among positive pairs (intra-class data points) and distinguish negative pairs (inter-class data points). SimCLR [4], for instance, has utilized instance-level comparisons unsupervised. However, this unsupervised approach raises concerns regarding its effectiveness , primarily because it limits the positive pairs to be transformed views of an image and treats all other samples in a mini-batch as negatives, potentially overlooking actual positive pairs. We hypothesis incorporating task-specific label information is thus crucial for accurately identifying all positive pairs, especially given the presence of labels in many downstream datasets.
@@
There is a growing trend of using task labels with contrastive learning to replace the standard use of CE [23]. A critical observation here is that many state-of-the-art methods, both in supervised [23, 15] and unsupervised [4, 48, 11, 57, 59] contrastive learning, overlook the strategic selection of negative samples. They fail to differentiate or prioritize these samples during selection or processing, thereby missing the benefits of leveraging “hard" negative samples, as highlighted in numerous studies [49, 25, 50, 5, 69, 26]. While contrastive learning mitigates the limitations of CE, it simultaneously introduces a challenge: a reliance on large batch sizes—such as 2048 or 4096 samples per batch—for superior performance compared to CE. This requirement is often impractical in budget hardware environments, particularly when using GPUs with less than 24 GB of memory. As a consequence, state-of-the-art methods such as SupCon [23] underperform compared to CE when using more commonly employed batch sizes, such as 64 or 128 samples per batch, which limits their application. Motivated by these successes and gaps in research, we pose the question: How can the performance of contrastive learning be improved to address the shortcomings of cross-entropy loss, while also mitigating the reliance on large batch sizes?
@@
  • State-of-the-Art Performance in Few-Shot Learning and Transfer Learning settings: CLCE significantly surpasses CE by an average of 2.74% in Top-1 accuracy across four few-shot learning datasets when using the BEiT-3 base model [56], with large gains observed in 1-shot learning scenarios. Additionally, in transfer learning settings, CLCE consistently outperforms other state-of-the-art methods across eight image datasets, setting a new state-of-the-art result for base models (88 million parameters) on ImageNet-1k [7].
  • Reduced Contrastive Learning’s Dependency on Large Batch Sizes: Empirical evidence shows that CLCE significantly outperforms both CE and previous state-of-the-art contrastive learning methods like SupCon [23] in commonly used batch sizes, such as 64. This is a size at which earlier state-of-the-art contrastive learning methods underperform. This advancement tackles a crucial bottleneck in contrastive learning, particularly in settings with limited resources. It positions CLCE as a viable, efficient alternative to conventional CE.
@@ 2.1 Limitations of Cross-Entropy loss
The cross-entropy loss (CE) has long been the default setting for many deep neural models due to its ability to optimize classification tasks effectively. However, recent research has revealed several inherent drawbacks [27, 3, 32]. Specifically, models trained with the CE tend to exhibit poor generalization capabilities. This vulnerability stems from the model having narrow decision margins in the feature space, making it more susceptible to errors introduced by noisy labels [39, 27] and adversarial examples [10, 39]. These deficiencies underscore the need for alternatives that offer better robustness and discrimination capabilities.
@@ 2.2 Contrastive Learning and Negative Mining
The exploration of negative samples, particularly hard negatives, in contrastive learning has emerged as a critical yet relatively underexplored area. While the significance of positive sample identification is well-established [60, 12, 68], recent studies have begun to unravel the intricate role of hard negatives. The potential of hard negative mining in latent spaces has been validated in numerous studies [50, 5, 69, 26, 58, 62, 13, 30, 34, 44]. These studies highlight the pivotal role of hard negatives in enhancing the discriminative capability of embeddings. In the contrastive learning domain, [6] tackled the challenge of discerning true negatives from a vast pool of candidates by approximating the true negative distribution. Later, [47] applied hard negative mining to unsupervised contrastive learning, resulting in a framework where only a single positive pair is utilized in each iteration of the loss calculation. However, these approaches still presents limitations, such as inaccurately identifying positive and negative samples and only using one positive pair, which harms the performance of contrastive learning. H-SCL [22] expand upon the concept of hard negative mining within a supervised framework. Although both our work and H-SCL utilize hard negative sampling, the methodologies for implementing sampling significantly differ between the two. Their approach employs a consistent threshold-based dot product for identifying “hard” samples. However, determining an appropriate threshold remains challenging, as it varies significantly across different datasets and even within individual mini-batches. In contrast, our CLCE method dynamically determines the weighting of each sample, proving to be significantly more effective than H-SCL. Moreover, their methodology does not tackle the dependency on large batch sizes, which is a critical limitation on performance and applicability.
@@ 3.1 CLCE
$\mathcal{L}_{\text{LACLN}}=\displaystyle\sum\limits_{x_{i}\in\mathcal{D}^{*}}-\log\frac{1}{\lvert\mathcal{D}^{*\textbf{+}}_{-x_{i}}\lvert}\frac{\displaystyle\sum\limits_{x_{p}\in\mathcal{D}^{*\textbf{+}}_{-x_{i}}}\exp(x_{i}\cdot x_{p}/\tau)}{\displaystyle\sum\limits_{x_{p}\in\mathcal{D}^{*\textbf{+}}_{-x_{i}}}\exp(x_{i}\cdot x_{p}/\tau)+\displaystyle\sum\limits_{x_{k}\in\mathcal{D}^{*\textbf{-}}_{-x_{i}}}\displaystyle\frac{\lvert\mathcal{D}^{*\textbf{-}}_{-x_{i}}\lvert}{\displaystyle\sum\limits_{x_{k}\in\mathcal{D}^{*\textbf{-}}_{-x_{i}}}{\exp(x_{i}\cdot x_{k}/\tau)}}\exp(x_{i}\cdot x_{k}/\tau)^{2}}$
@@ 3.3 Representation Learning Framework
  • Data Augmentation module, ${Aug(\cdot)}$: This component creates two different views of each sample $r$, denoted $\tilde{r}={Aug(r)}$. This means that every sample will have at least one similar sample (positive pair) in a batch during training.
  • Encoder Network, ${Enc(\cdot)}$: This network encodes the input data, $r$, into a representation vector, $x=Enc(r)$. Each of the two different views of the data is fed into the encoder separately.
  • Classification head, $Head(\cdot)$: This maps the representation vector, $x$, to probabilities of classes in the target task. The mapping primarily consists of a linear layer, and we utilize its output to calculate the cross-entropy loss.
Our CLCE approach (Eq. 3) can be applied using a wide range of encoders, such as BEiT-3 [56] or the ResNets [16] for image classification. Following the method in [4], every image in a batch is altered to produce two separate views (anchors). Views with the same label as the anchor are considered positive, while the rest are viewed as negative. The encoder output, represented by $x_{i}={Enc(r_{i})}$, is used to calculate the contrastive loss. In contrast, the output from the classification head, denoted as $z_{i}={Head(Enc(r_{i}))}$, is used for the CE. We have incorporated L2 normalization on encoder outputs, a strategy demonstrated to enhance performance significantly [52].
@@ 4 Evaluation
We employ representative models from two categories of architectures – BEiT-3/MAE/ViT base [56, 17, 9] (transformers based models), and ResNet-101 [16] (convolutional neural network). While new state-of-the-art models are continuously emerging (e.g. DINOv2 [41]), our focus is not on the specific choice of architecture. Instead, we aim to show that CLCE is model-agnostic by demonstrating performance gains with two very different and widely used architectures, as well as show it can be trained and deployed in hardware-constrained settings. Further implementation details and the complete code for all experiments are publicly available at https://github.com/longkukuhi/CLCE.
@@
[8] Transductive 76.58$\pm$0.68 85.79$\pm$0.50 43.16$\pm$0.59 57.57$\pm$0.55 65.73$\pm$0.68 78.40$\pm$0.52 73.34$\pm$0.71 85.50$\pm$0.50
[67] Meta-QDA 75.83$\pm$0.88 88.79$\pm$0.75 - - 67.83$\pm$0.64 84.28$\pm$0.69 74.33$\pm$0.65 89.56$\pm$0.79
[18] FewTRUE-ViT 76.10$\pm$0.88 86.14$\pm$0.64 46.20$\pm$0.79 63.14$\pm$0.73 68.02$\pm$0.88 84.51$\pm$0.53 72.96$\pm$0.92 87.79$\pm$0.67
[18] FewTRUE-Swin 77.76$\pm$0.81 88.90$\pm$0.59 47.68$\pm$0.78 63.81$\pm$0.75 72.40$\pm$0.78 86.38$\pm$0.49 76.32$\pm$0.87 89.96$\pm$0.55
[21] BAVARDAGE 82.68$\pm$0.25 89.97$\pm$0.18 52.60$\pm$0.32 65.35$\pm$0.25 77.85$\pm$0.28 88.02$\pm$0.14 79.38$\pm$0.29 88.04$\pm$0.18
@@ 4.1 Few-shot Learning
We evaluate our proposed CLCE in the few-shot learning setting. The experiments on few-shot learning aim to assess the quality of the learned representations. Specifically, each test run comprises 3,000 randomly sampled tasks, and we report median Top-1 accuracy with a 95% confidence interval across three runs, maintaining a consistent query shot count of 15. Four prominent benchmarks are used for evaluation: CIFAR-FS [2], FC100 [42], miniImageNet [54], and tieredImageNet [46]. We follow established splitting protocols for a fair comparison [2, 42, 45].
Tab. 1 shows the performance of BEiT-3 and ResNet-101 models under various methods, including CE, H-SCL [22], and the same weighted combination of CE and state-of-the-art supervised contrastive learning loss (SupCon) [23] as CLCE. The results reveal that our CLCE approach consistently improves classification accuracy over other methods, demonstrating superior generalization with limited training data for each class. Our CLCE enhances models’ performance on few-shot datasets, significantly outperforming both CE and CE+SuperCon (paired t-test, p $<$ 0.01). In the 1-shot learning context when compared to BEiT-3 trained with CE (BEiT-3-CE), the most remarkable improvement is seen on the FC100 dataset, with accuracy rising by $3.52\%$ through the use of CLCE (BEiT-3-CLCE). Indeed, across all datasets, BEiT-3-CLCE shows an average accuracy improvement of $2.7\%$. For 5-shot learning, the average improvements across the datasets are $1.4\%$ in accuracy for BEiT-3-CLCE, demonstrating CLCE’s effectiveness in scenarios with fewer positive samples per class and its ability to yield consistent and reliable results, evident in the tighter confidence intervals for Top-1 accuracy. As for ResNet-101, CLCE (ResNet-101-CLCE) demonstrates even more significant improvements over both CE and CE+SupCon. The enhancement is especially remarkable in the case of tieredImagenet, where ResNet-101-CLCE achieves increases of $16.68\%$ over ResNet-101-CE and $14.17\%$ over ResNet-101-CE+SupCon in 1-shot learning. For 5-shot learning, the improvements are $16.9\%$ and $13.36\%$, respectively. On average, ResNet-101-CLCE achieves a $9.82\%$ improvement in 1-shot and an $8.71\%$ improvement in 5-shot settings over the ResNet-101-CE. Lastly, H-SCL [22] underperforms compared to CE at a batch size of 128. This highlights contrastive learning’s limitation of needing very large batch sizes for better performance than CE, evident in ResNet-101 and BEiT-3 models.
Overall, the enhancement of our CLCE is particularly effective for few-shot scenarios, where limited labelled data requires the model to rely more on high-quality, discriminative representations. These outcomes underline the efficacy of our proposed CLCE approach and CLCE’s broad applicability across different model architectures for few-shot learning tasks.